BOND-STRETCH ISOMERISM IN SFCl AND OTHER SULFUR HALIDES

\underline{JEFF LEIDING}, DAVID E. WOON and THOM H. DUNNING, JR., Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana IL, 61801.

We have identified bond-stretch isomers in ^{3}A′′ excited states of SFCl and other triatomic sulfur halides using high-level ab initio techniques [RCCSD(T)/aug-cc-pVQZ]. These bond-stretch isomers are true minima on their potential energy surfaces, and they are manifested by alternation in bond type: There is a shorter (covalent) bond, and longer (recoupled pair) bond in all cases, while the bond angle is nearly constant for the isomers (∼90^{∘}). The SFCl isomers are ∼15 kcal/mol above a third nearly linear isomer on the PES for which both bonds are recoupled pair bonds. The recoupled pair bond is one in which the p^{2} pair of singlet coupled electrons on the sulfur has been decoupled so that one of the electrons is recoupled to form a bond with a halogen and is typically ∼0.3 Å longer than a covalent bond. The bond-stretch isomers are interesting because the potential in the direction of bond alternation is very flat, and the barrier between the isomers is very low (<1.0 kcal/mol for SFCl, for example). Harmonic frequencies of these minima and the transition states between them have been calculated. Potential energy surfaces are being calculated in order to obtain a numerical solution to the nuclear Schrödinger equation.